# LlamaRestTest: Effective REST API Testing with Small Language Models

### v1

@@ Abstract.
Modern web services rely heavily on REST APIs, typically documented using the OpenAPI specification. The widespread adoption of this standard has resulted in the development of many black-box testing tools that generate tests based on these specifications. Recent advancements in Natural Language Processing (NLP), particularly with Large Language Models (LLMs), have enhanced REST API testing by extracting actionable rules and generating input values from the human-readable portions of the specification. However, these advancements overlook the potential of continuously refining the identified rules and test inputs based on server responses. To address this limitation, we present LlamaRestTest, a novel approach that employs two custom LLMs to generate realistic test inputs and uncover parameter dependencies during the testing process by incorporating server responses. These LLMs are created by fine-tuning the Llama3-8b model, using mined datasets of REST API example values and inter-parameter dependencies. We evaluated LlamaRestTest on 12 real-world services (including popular services such as Spotify), comparing it against RESTGPT, a GPT-powered specification-enhancement tool, as well as several state-of-the-art REST API testing tools, including RESTler, MoRest, EvoMaster, and ARAT-RL. Our results demonstrate that fine-tuning enables smaller LLMs to outperform much larger models in detecting actionable rules and generating inputs for REST API testing. We also evaluated different tool configurations, ranging from the base Llama3-8B model to fine-tuned versions tailored for REST API testing, and explored multiple quantization techniques, including 2-bit, 4-bit, and 8-bit integer formats, to improve efficiency and performance. Our study shows that small language models can perform similar to, or better than, large language models in REST API testing, striking a balance between effectiveness and efficiency. Furthermore, LlamaRestTest outperforms state-of-the-art REST API testing tools in code coverage achieved and internal server errors identified, even when those tools utilize enhanced specifications generated by RESTGPT. Finally, through an ablation study, we show that each of the novel components of LlamaRestTest contributes to its overall performance.
@@
journal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Web services often rely on Representational State Transfer (REST) APIs for efficient communication and data exchange. REST, a paradigm established by Roy Fielding (Fielding, 2000), emphasizes scalability and flexibility, enabling seamless interactions between diverse software systems through standard web protocols (Pautasso et al., 2008). REST APIs are pivotal in modern software architecture, facilitating operations over the Internet without requiring client and server to know details about each other beyond the shared understanding communicated through the API itself (Richardson et al., 2013).
@@
Platforms like APIs Guru (APIs.guru, 2023) and Rapid API (R Software Inc., 2023) highlight the prevalence of REST APIs by hosting a vast collection of REST API specifications and emphasizing the critical role of clear, standardized documentation. REST APIs are typically documented using structured languages such as OpenAPI (OpenAPI, 2023), RAML (MuleSoft, LLC, a Salesforce company, 2020), and API Blueprint (API Blueprint, 2023). These specifications provide essential API details, such as available operations, data formats, and response types, offering a structured way of describing REST APIs, thus enhancing comprehensibility and compatibility with software tools. This structured approach to documentation sparked the development of many automated black-box testing tools for REST APIs (e.g., (Corradini et al., 2022; Atlidakis et al., 2019; Karlsson et al., 2020; Martin-Lopez et al., 2021; Karlsson et al., 2021; Hatfield-Dodds and Dygalo, 2022; Wu et al., 2022; Kim et al., 2023b; tcases, 2023; Liu et al., 2022)).
@@
However, despite advancements in REST API testing tools, significant challenges remain due to constraints and inter-parameter dependencies among API parameters (Kim et al., 2022; Martin-Lopez et al., 2019). Current REST API testing tools employ different techniques, such as breadth-first search (BFS), depth-first search (DFS), combinatorial testing, evolutionary algorithms, random-based approaches, and Q-Learning, to address these challenges. Among these tools, NLP2REST (Kim et al., 2023a) employs Natural Language Processing (NLP) to extract actionable REST API rules and converts them into machine-readable formats for REST API testing tools. Similarly, ARTE (Alonso et al., 2023) generates realistic parameter inputs by extracting key information from parameter names and descriptions and querying knowledge bases such as DBPedia (Bizer et al., 2009) for input values. RESTGPT (Kim et al., 2023c) uses Large Language Models (LLMs) to perform both actionable rule extraction and test input generation. Although these tools achieve considerable success in enhancing API specifications and improving testing effectiveness, they lack the ability to dynamically interact with testing tools during execution. This limitation prevents them from refining test inputs and rules based on real-time server feedback, leaving both critical inter-parameter dependencies and semantically valid value generation under-explored.
To address these limitations, we present LlamaRestTest, the first black-box testing approach that leverages the human-like reasoning capabilities of LLMs (OpenAI, 2023; Touvron et al., 2023) to take into account dynamic server feedback during testing. The core novelty of LlamaRestTest lies in its two components, LlamaREST-IPD and LlamaREST-EX, which are LLMs created by fine-tuning the Llama3-8b model (Meta, 2024b) for the tasks of identifying inter-parameter dependencies and generating input values for operation parameters, respectively. LlamaREST-IPD decides which parameters to select for each API, whereas LlamaREST-EX determines what values to assign to each parameter. LlamaREST-IPD is fine-tuned using a database used in a study by Martin-Lopez (Martin-Lopez et al., 2019), which includes a range of parameters and natural-language descriptions of their dependencies. We parse these parameters and their associated actionable inter-parameter dependency rules using RESTGPT (Kim et al., 2023c) and manually validate the rules to create the fine-tuning database. LlamaREST-EX is fine-tuned on a dataset of example values mined from over four thousand API specifications, containing more than 1.8 million parameters, from APIs-guru (APIs.guru, 2023). We collected the parameters which contain example values, and through fine-tuning, we enhance LlamaREST-EX’s ability to generate realistic testing inputs for REST APIs. Both these models are then quantized to reduce their sizes, resulting in smaller, faster, streamlined model versions. We integrated these LLMs with the ARAT-RL framework (Kim et al., 2023b), which is a reinforcement learning based REST API testing tool. When ARAT-RL struggles to process an API operation, LlamaREST-IPD and LlamaREST-EX utilize both parameter descriptions and server response messages to provide dependency rules and example values, enabling the generation of valid parameter sets and values for the parameters.
@@
To assess the performance of LlamaRestTest, we collected 12 real-world RESTful services used in recent studies (Kim et al., 2023a, b), including popular services such as FDIC and Spotify. We compared our tool with RESTGPT to evaluate the effectiveness of LlamaREST-IPD and LlamaREST-EX using various Llama3-8B configurations, including the vanilla Llama3-8B model, fine-tuned (FT), and fine-tuned with 2-bit, 4-bit, and 8-bit quantization. Additionally, we explored how fine-tuning and quantization impact overall testing performance. We then compared LlamaRestTest with four state-of-the-art REST testing tools recognized as top performers in recent research (Kim et al., 2022, 2023b; Liu et al., 2022; Zhang and Arcuri, 2023): RESTler (Atlidakis et al., 2019), EvoMaster (Arcuri, 2019), MoRest (Liu et al., 2022), and ARAT-RL (Kim et al., 2023b). We measured effectiveness using code coverage for open-source services, operation coverage for online services, and the ability to detect internal server errors across all services, as these are the most commonly used metrics in the field (Golmohammadi et al., 2023; Kim et al., 2022). Finally, we conducted an ablation study to assess how specification information, server response messages, parameter input generation, and parameter dependency identification each contribute to LlamaRestTest’s overall performance.
RESTGPT generated valid inputs for 68.82% of parameters, while base Llama3-8B achieved 22.94%. After fine-tuning, LlamaREST-EX improved to 72.44%. Quantized versions (8-bit, 4-bit, 2-bit) achieved 66.12%, 60.21%, and 29.12% success rates. For inter-parameter dependencies, RESTGPT detected 9 out of 17 cases, base Llama3-8B found 2, and LlamaREST-IPD identified 12. Quantized versions detected 1 (2-bit), 6 (4-bit), and 7 (8-bit). These results show that fine-tuning significantly improves accuracy, and 4-bit and 8-bit quantized versions maintain reasonable performance.
The FT with 8-bit quantization model achieved the highest coverage (55.8% method, 28.3% branch, and 55.3% line coverage), followed by the 4-bit and 2-bit models. The vanilla model had the lowest coverage: 49.2%, 23.2%, and 48.5%. In terms of processed operations, the FT with 8-bit model achieved the best result, with 37.5 operations, whereas the vanilla model processed 23. These results highlight the importance of balance between the models’ effectiveness and efficiency.
Compared to other state-of-the-art testing tools using RESTGPT’s enhanced specifications, LlamaRestTest achieved between 59% and 194% more branch coverage, between 22% and 65% more line coverage, and between 22% and 67% more method coverage than its competitors. LlamaRestTest also processed 37.5 operations per run, far surpassing EvoMaster (11.2), ARAT-RL (11.6), RESTler (10.6), and MoRest (9.9). Additionally, it identified 204 internal server errors, outperforming ARAT-RL (160), MoRest (140), and both RESTler and EvoMaster (130).
Our ablation study shows that the most significant impact was caused by removing LlamaREST-IPD, which reduced method, branch, and line coverage by 16.5%, 21.9%, and 15.2%, respectively. Server response removal also caused major drops (10.9%, 15.5%, and 11.4%), while removing parameter descriptions led to smaller declines (10.6%, 4.9%, and 8.1%). Lastly, removing LlamaREST-EX reduced coverage by 6.6%, 14.7%, and 5.4%. These results emphasize the importance of each component in the performance of our approach.
@@
  • LlamaRestTest, the first language model-based REST API testing approach that leverages server feedback along with fine-tuning and quantization techniques to improve both effectiveness and efficiency in REST API testing.
  • Empirical results that show the effectiveness of fine-tuning and quantization in REST API testing and show that LlamaRestTest outperforms the current state-of-the-art testing tools.
  • An artifact (Anonymous, 2023) that includes the LlamaRestTest tool, benchmarking services, and the database used for fine-tuning, along with an extensive set of empirical data.
@@ 2. Background and Motivating Example
/elements/area:
get:
summary: Area of OSM elements
operationId: area
parameters:
- name: bboxes
in: query
description: WGS84 coordinates in the following format: id1:lon1,lat1,lon2,lat2|id2:lon1,lat1,lon2,lat2|… OR lon1,lat1,lon2,lat2|lon1,lat1,lon2,lat2|…; no default value (one boundary parameter must be defined)
type: string
- name: bcircles
in: query
description: WGS84 coordinates + radius in meter in the following format: id1:lon,lat,r|id2:lon,lat,r|… OR lon,lat,r|lon,lat,r|…; no default value (one boundary parameter must be defined)
type: string
- name: bpolys
in: query
description: WGS84 coordinates given as a list of coordinate pairs (as for bboxes) or GeoJSON FeatureCollection. The first point has to be the same as the last point and MultiPolygons are only supported in GeoJSON; no default value (one boundary parameter must be defined)
type: string
- name: filter
in: query
description: Combines several attributive filters, e.g. OSM type, the geometry (simple feature) type, as well as the OSM tag; no default value
type: string
- name: format
in: query
description: Output format geojson (for /groupBy/boundary resources only), csv, or json;
type: string
- name: time
in: query
description: ISO-8601 conform timestring(s); default: latest timestamp within dataset
type: string
responses:
’200’:
description: OK
@@ 2.1. REST APIs and OpenAPI Specification
REST APIs are web APIs that adhere to RESTful principles, enabling data exchange between clients and servers using protocols like HTTP (Berners-Lee et al., 1996). Key concepts of REST include statelessness, where each request is independent; cacheability, allowing response storage; and a uniform interface that simplifies client-server interactions (Fielding, 2000; Richardson et al., 2013). Clients interact with web resources via HTTP requests, performing operations such as creating, reading, updating, or deleting data using methods like POST, GET, PUT, and DELETE. Each HTTP method paired with a resource endpoint defines an operation, with requests containing headers, payloads, and metadata. Responses include headers, bodies, and status codes indicating success (2xx), client errors (4xx), or internal server errors (500), forming the foundation of RESTful communication.
@@
The OpenAPI Specification (OAS) (OpenAPI, 2023), which evolved from Swagger in version 2 to the OpenAPI Specification in version 3, is a crucial standard in the realm of RESTful API design and documentation. As an industry-standard format, OAS delineates the structure, functionalities, and anticipated behaviors of APIs in a standardized and human-accessible way. Figure 1 presents a portion of the Ohsome API’s OpenAPI Specification. This example highlights the structured approach of OAS in defining API operations. For instance, it specifies a GET request at the /elements/area endpoint, aimed at calculating the area of OpenStreetMap (OSM) elements. This operation contains multiple query parameters, such as bboxes, bcircles, and bpolys, each requiring specific coordinate formats and types. Parameters like filter, format, and time are also included, offering additional functionality and flexibility in data querying. The API’s response to a successful query typically includes a 200 status code, indicating a standard successful operation.
2.2. REST API Testing and Tools for Enhancing OpenAPI Specifications (OAS)

### v2

@@ Abstract.
Modern web services rely heavily on REST APIs, typically documented using the OpenAPI specification. The widespread adoption of this standard has resulted in the development of many black-box testing tools that generate tests based on OpenAPI specifications. Although Large Language Models (LLMs) have shown promising test-generation abilities, their application to REST API testing remains mostly unexplored. We present LlamaRestTest, a novel approach that employs two custom LLMs—created by fine-tuning and quantizing the Llama3-8B model using mined datasets of REST API example values and inter-parameter dependencies—to generate realistic test inputs and uncover inter-parameter dependencies during the testing process by analyzing server responses. We evaluated LlamaRestTest on 12 real-world services (including popular services such as Spotify), comparing it against RESTGPT, a GPT-powered specification-enhancement tool, as well as several state-of-the-art REST API testing tools, including RESTler, MoRest, EvoMaster, and ARAT-RL. Our results demonstrate that fine-tuning enables smaller models to outperform much larger models in detecting actionable parameter-dependency rules and generating valid inputs for REST API testing. We also evaluated different tool configurations, ranging from the base Llama3-8B model to fine-tuned versions, and explored multiple quantization techniques, including 2-bit, 4-bit, and 8-bit integer formats. Our study shows that small language models can perform as well as, or better than, large language models in REST API testing, balancing effectiveness and efficiency. Furthermore, LlamaRestTest outperforms state-of-the-art REST API testing tools in code coverage achieved and internal server errors identified, even when those tools use RESTGPT-enhanced specifications. Finally, through an ablation study, we show that each component of LlamaRestTest contributes to its overall performance.
@@
copyright: ccdoi: 10.1145/3715737journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE022publicationmonth: 7ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Web services often rely on Representational State Transfer (REST) APIs for efficient communication and data exchange. REST, a paradigm established by Roy Fielding (Fielding, 2000), emphasizes scalability and flexibility, enabling seamless interactions between diverse software systems through standard web protocols (Pautasso et al., 2008). REST APIs are pivotal in modern software architecture, facilitating operations over the internet without requiring the client and the server to know details about each other beyond the shared understanding communicated through the API itself (Richardson et al., 2013). Given their widespread adoption and critical role in modern web services, ensuring proper documentation and testing of REST APIs has become increasingly important.
@@
Platforms like APIs Guru (APIs.guru, 2023) and Rapid API (R Software Inc., 2023) highlight the prevalence of REST APIs by hosting a vast collection of REST API specifications and highlighting the critical role of clear and standardized documentation. REST APIs are typically documented using structured languages such as OpenAPI Specification (OAS) (OpenAPI, 2023) (formerly known as Swagger (Software, 2023)), RAML (MuleSoft, LLC, a Salesforce company, 2020), and API Blueprint (API Blueprint, 2023). These specifications provide essential API details, such as available operations, data formats, and response types, offering a structured way of describing REST APIs.
This structured API documentation has led to the development of many automated black-box testing tools for REST APIs (e.g., (Corradini et al., 2022; Atlidakis et al., 2019; Karlsson et al., 2020; Martin-Lopez et al., 2021; Karlsson et al., 2021; Hatfield-Dodds and Dygalo, 2022; Wu et al., 2022; Kim et al., 2024a; Kimbrough, 2023; Liu et al., 2022; Kim et al., 2025; Corradini et al., 2024; Lin et al., 2023)) that take API specifications as input and drive testing from the specifications. Current REST API testing tools employ different techniques, including stateful fuzzing, combinatorial testing, evolutionary algorithms, deep learning, and Q-Learning, to perform effective testing. However, despite advancements in REST API testing, significant challenges remain due to complex input constraints and inter-parameter dependencies among API parameters of real-world RESTful services (Kim et al., 2022; Martin-Lopez et al., 2019).
@@
To address these challenges, researchers have turned to recent advances in the field of natural language processing (NLP), leading to several promising approaches for REST API testing. For example, NLP2REST (Kim et al., 2023) employs NLP techniques to extract constraints and dependencies from OAS and convert them into OAS keywords (e.g., minimum, maximum, pattern) that testing tools can use. ARTE (Alonso et al., 2023) generates realistic parameter inputs by extracting key information from parameter names and descriptions and querying knowledge bases, such as DBPedia (Bizer et al., 2009). RESTGPT (Kim et al., 2024b) uses Large Language Models (LLMs) to perform rule extraction and parameter input generation. These tools have improved REST API testing effectiveness by extracting additional machine-readable (structured) information from the human-readable portions of API specifications. However, they lack the ability to dynamically interact with testing tools during execution. This limitation prevents them from refining test inputs and rules based on real-time server feedback.
To address these limitations, we present LlamaRestTest, the first black-box testing approach that leverages Language Models (LMs) to incorporate dynamic server feedback during REST API testing. LlamaRestTest consists of two specialized LMs created by fine-tuning the Llama3-8b model (Meta, 2024b): LlamaREST-IPD for identifying inter-parameter dependencies, and LlamaREST-EX for generating appropriate input values. These models were fine-tuned using a comprehensive dataset combining Martin-Lopez’s dependency database (Martin-Lopez et al., 2019) and over 1.8 million API operation parameters mined from APIs-guru (APIs.guru, 2023). We also created quantized versions of these models to investigate their efficiency-accuracy tradeoff when using lower precision. LlamaRestTest further leverages these models together with reinforcement learning for more effective API exploration. Specifically, we integrated these models with ARAT-RL (Kim et al., 2024a) [^1] We chose ARAT-RL as its exploration strategy has been shown to be effective among the recent tools (Kim et al., 2024a)., a reinforcement learning-based testing approach, enabling dynamic analysis of parameter descriptions and server responses to generate valid parameter combinations and input values during testing.
@@
To assess the performance of LlamaRestTest, we collected 12 real-world RESTful services, previously evaluated in (Kim et al., 2023, 2024a), including popular services such as Spotify. We compared our tool with RESTGPT to evaluate the effectiveness of LlamaREST-IPD and LlamaREST-EX in identifying inter-parameter dependencies and computing valid input values, respectively, using various configurations, including the base Llama3-8B model, the fine-tuned model, and the fine-tuned model with 2-bit, 4-bit, and 8-bit quantization. Additionally, we investigated how fine-tuning and quantization impact overall testing performance. We then compared LlamaRestTest with four state-of-the-art REST testing tools: RESTler (Atlidakis et al., 2019), EvoMaster (Arcuri, 2019), MoRest (Liu et al., 2022), and ARAT-RL (Kim et al., 2024a). We measured effectiveness using three key metrics widely adopted in API testing research (Golmohammadi et al., 2023; Kim et al., 2022): code coverage (for open-source services), operation coverage (for online services), and the ability to detect internal server errors (for all services). Finally, we conducted an ablation study to assess how specification information, server response messages, parameter input generation, and parameter dependency identification each contribute to LlamaRestTest’s overall performance.
Our empirical evaluation demonstrated LlamaRestTest’s effectiveness across multiple dimensions. In parameter handling, our fine-tuned LlamaREST-EX model achieved a 72.44% success rate in generating valid inputs, outperforming the base Llama3-8B model (22.94%) while also being more effective than RESTGPT (68.82%). For inter-parameter dependencies, LlamaREST-IPD successfully identified 12 out of the 17 dependencies, whereas RESTGPT and the base Llama model identified nine and two dependencies, respectively. In terms of code coverage, the fine-tuned model with 8-bit quantization achieved the best balance, reaching the highest average coverage (55.8% method, 28.3% branch, and 55.3% line). This model processed 37.5 operations per run—14.5 more operations processed per run than the base Llama model (which processed 23 operations). Fine-tuning alone provided modest benefits (50.1% method, 23.5% branch, 49.9% line coverage, and 24.5 operations processed), showing improvements of 0.9 percentage points, 0.3 percentage points, 1.4 percentage points, and 1.5 additional operations respectively over the base model (49.2%, 23.2%, 48.5%, and 23 operations). The 4-bit quantized model maintained strong performance (54.0% method, 27.9% branch, 53.5% line coverage, and 34.4 operations processed), whereas the 2-bit model showed lower but still meaningful improvements (52.9% method, 25.7% branch, 52.6% line coverage, and 25.4 operations processed) compared to both the base and the fine-tuned model, highlighting the benefit of quantization in REST API testing.
When compared against state-of-the-art testing tools (EvoMaster, ARAT-RL, RESTler, and MoRest), LlamaRestTest demonstrated substantial improvements in coverage (by 10.0–22.7 percentage points for method coverage, 9.2–18.7 percentage points for branch coverage, and 10.0–22.1 percentage points for line coverage) while identifying more internal server errors (204 compared to 130–160). Our ablation study revealed that removing LlamaREST-IPD had the most significant impact, reducing coverage by 6.2–9.2 percentage points, followed by server response analysis (4.4–6.3 percentage points reduction), parameter descriptions (1.4–5.9 percentage points reduction), and LlamaREST-EX (3.0–4.5 percentage points reduction).
@@
  • LlamaRestTest, the first LM-based REST API testing approach that leverages server feedback along with fine-tuning and quantization techniques to improve both effectiveness and efficiency of REST API testing.
  • Empirical results that demonstrate the effectiveness of fine-tuning and quantization in REST API testing and show that LlamaRestTest outperforms other state-of-the-art testing tools.
  • An artifact (Kim, 2023) that includes the LlamaRestTest tool, benchmarking services, and the data used for fine-tuning, along with our empirical study results.
@@ 2. Background and Motivating Example
[…]
OpenAPI specification showing the /elements/area endpoint of Ohsome API, including parameters such as bboxes, bcircles, bpolys, filter, format, and time, along with their descriptions and types. /elements/area: get: summary: Area of OSM elements operationId: area parameters: - name: bboxes in: query description: WGS84 coordinates in the following format: id1:lon1,lat1,lon2,lat2|id2:lon1,lat1,lon2,lat2|… OR lon1,lat1,lon2,lat2|lon1,lat1,lon2,lat2|…; no default value (one boundary parameter must be defined) type: string - name: bcircles in: query description: WGS84 coordinates + radius in meter in the following format: id1:lon,lat,r|id2:lon,lat,r|… OR lon,lat,r|lon,lat,r|…; no default value (one boundary parameter must be defined) type: string - name: bpolys in: query description: WGS84 coordinates given as a list of coordinate pairs (as for bboxes) or GeoJSON FeatureCollection. The first point has to be the same as the last point and MultiPolygons are only supported in GeoJSON; no default value (one boundary parameter must be defined) type: string - name: filter in: query description: Combines several attributive filters, e.g. OSM type, the geometry (simple feature) type, as well as the OSM tag; no default value type: string - name: format in: query description: Output format geojson (for /groupBy/boundary resources only), csv, or json; type: string - name: time in: query description: ISO-8601 conform timestring(s); default: latest timestamp within dataset type: string responses: ’200’: description: OK
@@ 2.1. REST APIs and OpenAPI Specification
REST APIs are web APIs that adhere to RESTful principles, enabling data exchange between clients and servers using protocols like HTTP (Berners-Lee et al., 1996). Key concepts of REST include statelessness, where each request is independent; cacheability, allowing response storage; and a uniform interface that simplifies client-server interactions (Fielding, 2000; Richardson et al., 2013). Clients interact with web resources via HTTP requests, performing operations such as creating, reading, updating, or deleting data using methods like POST, GET, PUT, and DELETE. Each HTTP method paired with a resource endpoint defines an operation, with requests containing headers, payloads, and metadata. Responses consist of headers, bodies, and status codes indicating success (2xx), client errors (4xx), or server errors (500).
@@
To effectively document and standardize these REST API interactions, the software industry has widely adopted formal specification formats. OAS (OpenAPI, 2023) is a crucial standard in the realm of RESTful API design and documentation. As an industry-standard format, it delineates the structure, functionalities, and anticipated behaviors of APIs in a standardized and human-accessible way. Figure 1 presents a portion of the Ohsome API’s OAS and illustrates the structured approach of defining API operations in OAS. For instance, it specifies a GET request at the /elements/area endpoint, aimed at calculating the area of OpenStreetMap (OSM) elements. This operation contains multiple query parameters, such as bboxes, bcircles, and bpolys, each requiring specific coordinate formats and types. Parameters like filter, format, and time are also included, offering additional functionality and flexibility in data querying. The API’s response to a successful query typically includes a 2xx status code, indicating the success of the operation.
2.2. REST API Testing and Tools for Enhancing OpenAPI Specifications